Grade 11Physical Chemistry

1g mixture of Na2CO3 and MaHCO3 is treated with excess of dilute HCl. The volume of CO2 evolved was 240cc at 27°C and 778.6 mm of Hg pressure. Calculate the percentage composition of the mixture.

Debarshi Das
9 Years agoGrade 11

1 Answer

Vikas TU
9 Years ago
From ideal gas eqn. we get,
PV = nRT
n = (778.6*240)/(760*1000*300*0.0821) = 0.00998 moles of CO2.
Now let x be the weight of Na2CO3 present in the mixture.
Then the reaction is:
Na2CO3   +     2HCl    ==========>>>>>>.             2NaCl       +          CO2         +      H2O
moles of Na2CO3 => moles of CO2 =>  0.00998
mass of Na2CO3 => 0.00998*106 = 1.055 gms which is exceeding th 1 gm and therfore cannot be possible.
U need toc heck the qstn. or the reaction would not feasible.